OpenStudy (anonymous):

(3a^2-5a+2)(2a^2-3a+4)... I'm stuck at multiplying 3a^2 and 2a^2

OpenStudy (anonymous):

okay well, since these have the same base and exponet and we have to do is multiply the coeffiecnets, 3*2=6, and add the expoents (expoent rule a^m*a^n=a^m+n), thus we get 6a^4

OpenStudy (anonymous):

ok that's what i thought but wasn't entirely sure

OpenStudy (anonymous):

so before collecting like terms it would be 6a^4-15a+8?

OpenStudy (anonymous):

well, i havent done it, but that looks a little off....

OpenStudy (anonymous):

reason being that to expand this equation you would need to multiply all the terms in the second bracket by each term in the first brackket. Then collect like terms.

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

well, first you multip all the terms in the second bracket by the first term in the first brackket. So you multiply (2a^2-3a+4), by 3a^2.

OpenStudy (anonymous):

The you multiply (2a^2-3a+4) by -5a

OpenStudy (anonymous):

i thought i was finally starting to get this but apparently not

OpenStudy (anonymous):

this would be easier without all the letters and exponents

OpenStudy (anonymous):

it is just a process of repitition.

OpenStudy (anonymous):

Then multiply (2a^2-3a+4) by 2

OpenStudy (anonymous):

I will do it for you okay

OpenStudy (anonymous):

(3a^2-5a+2)(2a^2-3a+4) Step 1.6a^4-9a^3+12a^2 (here i multiplied all the terms in the second bracket by 3a^2) Step 2: -10a^3+15a^2-20a( here i multiplied all the terms in the second brackket by -5a) Step 3: 4a^2-6a+8( here i multiplied all the terms in the second brackket by 2) Step 4: 6a^4-9a^3+12a^2-10a^3+15a^2-20a+4a^2-6a+8( i gathered all the products i got in step 1-3) Step 5: 6a^4-19a^3+31a^2-26a+8( gather like terms together)

OpenStudy (anonymous):

thank you so much, I have it written down step by step so I can use it for later reference.